1.     Assume the following for price discriminating monopolist aimed at maximizing profit. Total demand for the product of the monopolist is Q = 50-5P (P = 100-2Q) (10 Marks)

·        Demand in Market one is Q1= 32-0.4P1(P1= 80-2.5Q1)

·        Demand in Market two is Q2= 18-0.1P2(P2= 180-10Q2)

·        Cost function is C = 50+40Q (where Q = Q1+ Q2)

a.      Find equilibrium quantities (Q1and Q2),

b.     Find equilibrium prices (P1and P2),

c.      Calculate profit (π),

d.     Calculate and Interpret elasticities (ε1 and ε2)

 



1
Expert's answer
2021-07-06T08:31:24-0400

a.

Total revenue in market 1

"= 80Q_1 - 2.5 Q_1^2"

Marginal revenue in market 1

"=\frac{ ∆TR}{ ∆ Q_1}\\

 = 80 - 5 Q_1 \\

MC = 40 \\

80 - 5 Q_1 = 40 \\

40 = 5 Q_1 \\

Q_1 = 8"


Total revenue 2

"= P \u00d7 Q = 180Q_2 - 10 Q_2^2"

Marginal revenue 2

 "= 180 - 20 Q_2 \\

MC = 40 \\

Equating \space MR = MC \\

180 - 20 Q_2 = 40 \\

140 = 20 Q2 

Q_2 = 7 \\"


b.

"P_1 = 80 - 2.5 (8) = 60\\

P_2 = 180-10Q_2 = 110"


c.

"Profit = TR_ 1 + TR_ 2 - TC \\

 = 8(60) + 7(110) - 50 - 40(15) \\

 = 480 + 770 - 50 - 600\\ 

 = 600"


d.

"Q_1= 32-0.4P_1 \\

\frac{∆Q_1}{∆P_1 }= - 0.4 \\

e = \frac{∆Q_1 }{ ∆ P_1} \times \frac{ P
_1 }{Q _1}\\

e = -0.4 \times (\frac{60 






}{ 8}) = - 3\\

Q_2= 18-0.1P_2 \\

\frac{∆Q2}{ ∆ P_2 }= -0.1 \\

e =\frac{ ∆Q2 }{∆ P_2 }\times \frac{P_2 }{Q _2}= -0.1(\frac{110}{7}) = -1.57"


Good 2 is less price elastic. 

A higher price is charged to the low elasticity segment and a lower price is charged to the high elasticity segment.